### v1

@@ Abstract
Class Incremental Learning (CIL) constitutes a pivotal subfield within continual learning, aimed at enabling models to progressively learn new classification tasks while retaining knowledge obtained from prior tasks. Although previous studies have predominantly focused on backward compatible approaches to mitigate catastrophic forgetting, recent investigations have introduced forward compatible methods to enhance performance on novel tasks and complement existing backward compatible methods. In this study, we introduce an effective-Rank based Feature Richness enhancement (RFR) method, designed for improving forward compatibility. Specifically, this method increases the effective rank of representations during the base session, thereby facilitating the incorporation of more informative features pertinent to unseen novel tasks. Consequently, RFR achieves dual objectives in backward and forward compatibility: minimizing feature extractor modifications and enhancing novel task performance, respectively. To validate the efficacy of our approach, we establish a theoretical connection between effective rank and the Shannon entropy of representations. Subsequently, we conduct comprehensive experiments by integrating RFR into eleven well-known CIL methods. Our results demonstrate the effectiveness of our approach in enhancing novel-task performance while mitigating catastrophic forgetting. Furthermore, our method notably improves the average incremental accuracy across all eleven cases examined.
@@ 3.1 Effective rank
Figure 2: Rank vs. feature richness. ResNet-18 was trained using ImageNet-100 dataset. (a) Supervised learning – as we include more classes in the training that starts from scratch, both $\mathsf{trank}$ and $\mathsf{erank}$ increase. Algebraic rank remains at the maximum value. (b) Unsupervised learning with SimCLR loss – as the unsupervised representation learning proceeds, both $\mathsf{trank}$ and $\mathsf{erank}$ increase. Algebraic rank remains at the maximum value.
@@ 4 Experiments
We provide an overview of our experimental settings in Section 4.1. Subsequently, we demonstrate the efficacy of our method for enhancing novel task performance in Section 4.2 and for mitigating catastrophic forgetting in Section 4.3. In Section 4.4, we demonstrate that our method can improve performance of eleven well-known backward-compatible methods. In Section 4.5, further analysis is provided. In Section 4.6, an ablation study is presented. In Section 4.2 and Section 4.3, we focus our analysis on the UCIR method, given its prevalent adoption across recent state-of-the-art methodologies.
@@
(a) S=10
(b) S=5
(c) S=2
4.1 Settings
Datasets:
We employ CIFAR-100 and ImageNet-100, two widely adopted benchmark datasets in CIL. To maintain consistency with prior studies, we follow the standard class orderings proposed in (Rebuffi et al., 2017) for all evaluations except for the evaluation of PODNet (Douillard et al., 2020) and IL2A (Zhu et al., 2021), for which we utilize the class orderings defined in PODNet and IL2A, respectively. To evaluate the CIL methods, all classes in datasets are divided into multiple tasks. The initial 50 classes are designated for the base task, while the remaining classes are split into novel tasks, where the size of each split is either 10, 5, or 2. In this study, we denote split sizes of 10, 5, and 2 as S=10, S=5, and S=2, respectively.
Implementation Details:
In this study, ResNet-18 (He et al., 2016) is employed as the base model to investigate. To reproduce the results of iCaRL (Rebuffi et al., 2017), LwF (Li and Hoiem, 2017), SI (Zenke et al., 2017), EEIL (Castro et al., 2018), MAS (Aljundi et al., 2018), RWalk (Chaudhry et al., 2018), BiC (Wu et al., 2019), IL2M (Belouadah and Popescu, 2019), and UCIR (Hou et al., 2019), we utilize the open-source codebase provided by FACIL (Masana et al., 2022). For PODNet (Douillard et al., 2020), IL2A (Zhu et al., 2021), and AFC (Kang et al., 2022), we employ their own open-source codebases. To incorporate RFR into the aforementioned methods, we modify their respective loss functions during the training of the base session, as specified by Eq. (6). Meanwhile, the remaining configurations, such as the loss functions applied during the training of novel sessions and the default hyper-parameter settings, remain unchanged. To evaluate the performance of the models, we adopt the standard metric of average incremental accuracy (AIC) proposed in (Rebuffi et al., 2017).
4.2 Superiority in forward compatibility
4.3 Mitigation in catastrophic forgetting
@@
To analyze improvement in on catastrophic forgetting, we conduct three investigations. First, we examine the influence of our method on the weight changes of the feature extractor. As novel sessions progress, the $L_{2}$-weight distance between the feature extractor learned in the base session and that in the novel session increases, as demonstrated in Figure 10. However, with the integration of RFR, the weight distances are significantly reduced across all sessions. The average reductions in weight distances are 8.70, 8.13, and 6.79 for split sizes of 10, 5, and 2, respectively. Additionally, Figure 10 reveals that the weight distance with the feature extractor from the immediate previous session also increases less when our method is employed, showing average reduction of 6.55, 4.95, and 2.99 for split sizes of 10, 5, and 2, respectively.
@@ 4.4 Consistent improvements over eleven existing methods
iCaRL (Rebuffi et al., 2017) 49.83±${}_{\pm 2.33}$ 46.86±${}_{\pm 2.38}$ 44.69±${}_{\pm 1.97}$ 53.77±${}_{\pm 1.68}$ 48.41±${}_{\pm 8.70}$ 49.78±${}_{\pm 3.65}$
  with RFR 50.56±${}_{\pm 1.47}$ 48.48±${}_{\pm 1.17}$ 44.99±${}_{\pm 0.97}$ 57.17±${}_{\pm 1.66}$ 55.62±${}_{\pm 0.71}$ 54.40±${}_{\pm 1.63}$
@@
EEIL (Castro et al., 2018) 37.81±${}_{\pm 2.55}$ 24.48±${}_{\pm 1.18}$ 32.37±${}_{\pm 2.52}$ 52.04±${}_{\pm 1.90}$ 48.65±${}_{\pm 2.34}$ 43.95±${}_{\pm 1.65}$
  with RFR 39.52±${}_{\pm 0.38}$ 25.79±${}_{\pm 1.18}$ 33.46±${}_{\pm 0.69}$ 54.74±${}_{\pm 0.91}$ 51.64±${}_{\pm 0.80}$ 47.89±${}_{\pm 0.60}$
@@
RWalk (Chaudhry et al., 2018) 35.75±${}_{\pm 1.63}$ 23.34±${}_{\pm 2.06}$ 27.02±${}_{\pm 5.36}$ 41.08±${}_{\pm 1.91}$ 21.44±${}_{\pm 5.69}$ 20.81±${}_{\pm 2.12}$
  with RFR 39.00±${}_{\pm 1.15}$ 24.44±${}_{\pm 1.06}$ 32.85±${}_{\pm 1.40}$ 44.83±${}_{\pm 1.87}$ 35.42±${}_{\pm 1.71}$ 22.68±${}_{\pm 4.70}$
@@
IL2M (Belouadah and Popescu, 2019) 40.85±${}_{\pm 2.22}$ 24.92±${}_{\pm 1.12}$ 33.41±${}_{\pm 2.12}$ 56.78±${}_{\pm 1.34}$ 52.35±${}_{\pm 2.16}$ 47.10±${}_{\pm 1.85}$
  with RFR 42.54±${}_{\pm 0.49}$ 25.19±${}_{\pm 0.41}$ 35.20±${}_{\pm 0.78}$ 58.40±${}_{\pm 1.04}$ 55.25±${}_{\pm 0.78}$ 50.09±${}_{\pm 0.83}$
@@
PODNet(NME) (Douillard et al., 2020) 68.47±${}_{\pm 1.27}$ 67.09±${}_{\pm 1.19}$ 65.16±${}_{\pm 1.02}$ 60.42±${}_{\pm 0.53}$ 50.93±${}_{\pm 0.73}$ 36.36±${}_{\pm 0.30}$
  with RFR 69.51±${}_{\pm 0.56}$ 67.66±${}_{\pm 0.63}$ 65.59±${}_{\pm 0.48}$ 64.08±${}_{\pm 0.45}$ 56.34±${}_{\pm 0.75}$ 41.03±${}_{\pm 2.15}$
@@
To demonstrate the efficacy of RFR in enhancing performance, we conduct comprehensive experiments over eleven well-known existing works. The results presented in Table 1 and Table 2 provide compelling evidence of the significant and consistent performance improvements achieved through the integration of RFR into the previous works. Specifically, we observe notable average performance improvements for the state-of-the-art works, including a 3.74% average improvement for UCIR, a 2.63% average improvement for PODNet (NME), a 1.24% average improvement for PODNet (CNN), and a 0.36% average improvement for AFC. Particularly, a remarkable performance improvement is observed in the case of UCIR for CIFAR-100 and its further analysis is shown in Figure 12. The achieved improvements are 3.14% increase for S=10, 5.59% for S=5, and 8.49% for S=2.
@@
ResNet-50 UCIR 67.10±${}_{\pm 0.20}$ 61.42±${}_{\pm 0.61}$ 51.62±${}_{\pm 0.14}$
  with RFR 68.57±${}_{\pm 0.12}$ 64.81±${}_{\pm 0.15}$ 57.71±${}_{\pm 0.31}$
@@
ViT-Small UCIR 50.45±${}_{\pm 0.35}$ 49.03±${}_{\pm 0.27}$ 46.24±${}_{\pm 0.79}$
  with RFR 51.86±${}_{\pm 0.35}$ 50.50±${}_{\pm 0.47}$ 47.96±${}_{\pm 1.03}$
@@ 4.5 Further analysis
In addition to evaluating RFR’s performance under the widely used benchmark settings, we extend our evaluation to a larger-scale dataset. As shown in Table 4, RFR continues to demonstrate its effectiveness for ImageNet-1000 dataset. Furthermore, we analyze RFR’s performance with larger-scale models. The results presented in Table 5 underscore RFR’s efficacy across a range of larger-scale models, including ResNet-50 and ViT-Small (Lee et al., 2021).
@@ 4.6 Ablation study
B=10 60.31±${}_{\pm 0.22}$ 60.41±${}_{\pm 0.34}$ +0.10
B=20 63.16±${}_{\pm 0.34}$ 65.11±${}_{\pm 0.29}$ +1.95
B=30 65.17±${}_{\pm 0.25}$ 66.98±${}_{\pm 0.21}$ +1.81
B=40 67.22±${}_{\pm 0.25}$ 69.53±${}_{\pm 0.12}$ +2.31
B=50 66.30±${}_{\pm 0.36}$ 69.45±${}_{\pm 0.29}$ +3.15
B=60 70.44±${}_{\pm 0.13}$ 73.06±${}_{\pm 0.36}$ +2.62
B=70 71.57±${}_{\pm 0.39}$ 74.20±${}_{\pm 0.20}$ +2.63
B=80 73.07±${}_{\pm 0.18}$ 75.95±${}_{\pm 0.09}$ +2.88
B=90 73.41±${}_{\pm 0.43}$ 75.98±${}_{\pm 0.19}$ +2.57
@@ 5.1 Comparison with CwD
Baseline 66.30±${}_{\pm 0.36}$ ${}_{\pm 0.56}$ 52.74±${}_{\pm 0.72}$ 66.11±${}_{\pm 0.63}$ 63.49±${}_{\pm 0.45}$ 59.61±${}_{\pm 0.53}$
  with CwD 67.06±${}_{\pm 0.12}$ 62.71±${}_{\pm 0.36}$ 56.39±${}_{\pm 0.30}$ 66.01±${}_{\pm 0.75}$ 63.63±${}_{\pm 1.02}$ 60.14±${}_{\pm 1.16}$
  with RFR 69.45±${}_{\pm 0.29}$ 66.16±${}_{\pm 0.10}$ 61.23±${}_{\pm 0.13}$ 67.13±${}_{\pm 0.86}$ 64.72±${}_{\pm 1.01}$ 61.49±${}_{\pm 0.71}$
@@ 5.2 Increasing representation rank during novel sessions
RFR increases effective rank during the base session only. It is also possible to increase effective rank during novel sessions with the goal of acquiring additional features that may be useful in the subsequent novel sessions. Such a strategy, however, can also intensify the modifications of the learned model from the previous sessions. To investigate the overall effect, we have conducted an experiment and the results are shown in Table 8. While RFR demonstrates effectiveness in enhancing feature richness during the base session, its influence on overall performance is marginal when also applied to novel sessions. Therefore, we have chosen to apply effective rank regularization only during the base session.
@@
BiC 58.06±${}_{\pm 0.73}$ 58.29±${}_{\pm 0.38}$ +0.23 48.02±${}_{\pm 2.33}$ 46.60±${}_{\pm 1.91}$1.41 27.74±${}_{\pm 3.11}$ 27.64±${}_{\pm 1.83}$0.10
EEIL 39.52±${}_{\pm 0.38}$ 39.44±${}_{\pm 0.63}$0.08 25.79±${}_{\pm 1.18}$ 25.62±${}_{\pm 1.10}$0.17 33.46±${}_{\pm 0.69}$ 33.16±${}_{\pm 1.07}$0.29
iCaRL 50.56±${}_{\pm 1.47}$ 50.50±${}_{\pm 2.04}$0.06 48.48±${}_{\pm 1.17}$ 48.23±${}_{\pm 1.14}$0.25 44.99±${}_{\pm 0.97}$ 44.81±${}_{\pm 1.18}$0.18
IL2M 42.54±${}_{\pm 0.49}$ 42.76±${}_{\pm 0.83}$ +0.21 25.19±${}_{\pm 0.41}$ 26.22±${}_{\pm 1.77}$ +1.03 35.20±${}_{\pm 0.78}$ 34.87±${}_{\pm 0.67}$0.32
LwF 40.98±${}_{\pm 0.34}$ 40.69±${}_{\pm 0.81}$0.28 28.82±${}_{\pm 4.17}$ 26.70±${}_{\pm 0.39}$2.12 34.46±${}_{\pm 0.63}$ 34.63±${}_{\pm 0.75}$ +0.17
MAS 39.54±${}_{\pm 0.71}$ 39.77±${}_{\pm 0.66}$ +0.23 28.80±${}_{\pm 4.52}$ 29.03±${}_{\pm 3.50}$ +0.23 33.36±${}_{\pm 0.52}$ 33.91±${}_{\pm 0.91}$ +0.55
SI 39.41±${}_{\pm 0.60}$ 39.51±${}_{\pm 0.58}$ +0.10 24.02±${}_{\pm 0.38}$ 23.89±${}_{\pm 0.19}$0.13 33.48±${}_{\pm 0.80}$ 33.60±${}_{\pm 0.52}$ +0.12
RWalk 39.00±${}_{\pm 1.15}$ 38.91±${}_{\pm 1.39}$0.09 24.44±${}_{\pm 1.06}$ 25.96±${}_{\pm 3.06}$ +1.53 32.85±${}_{\pm 1.40}$ 33.04±${}_{\pm 1.20}$ +0.19
UCIR 69.45±${}_{\pm 0.29}$ 69.57±${}_{\pm 0.23}$ +0.12 66.16±${}_{\pm 0.10}$ 66.55±${}_{\pm 0.16}$ +0.39 61.23±${}_{\pm 0.13}$ 61.27±${}_{\pm 0.29}$ +0.04

### v2

@@ Abstract
Class Incremental Learning (CIL) constitutes a pivotal subfield within continual learning, aimed at enabling models to progressively learn new classification tasks while retaining knowledge obtained from prior tasks. Although previous studies have predominantly focused on backward compatible approaches to mitigate catastrophic forgetting, recent investigations have introduced forward compatible methods to enhance performance on novel tasks and complement existing backward compatible methods. In this study, we introduce effective-Rank based Feature Richness enhancement (RFR) method that is designed for improving forward compatibility. Specifically, this method increases the effective rank of representations during the base session, thereby facilitating the incorporation of more informative features pertinent to unseen novel tasks. Consequently, RFR achieves dual objectives in backward and forward compatibility: minimizing feature extractor modifications and enhancing novel task performance, respectively. To validate the efficacy of our approach, we establish a theoretical connection between effective rank and the Shannon entropy of representations. Subsequently, we conduct comprehensive experiments by integrating RFR into eleven well-known CIL methods. Our results demonstrate the effectiveness of our approach in enhancing novel-task performance while mitigating catastrophic forgetting. Furthermore, our method notably improves the average incremental accuracy across all eleven cases examined.
@@ 1 Introduction
The main contributions can be summarized as follows: (1) While previous methods focus solely on either forward compatibility or backward compatibility, this study proposes the novel method RFR, which enhances both forward and backward compatibility simultaneously. (2) This study provides both theoretical and empirical evidence supporting the superiority of the method.
@@ 3.1 Effective rank
Figure 2: Rank vs. feature richness. ResNet-18 was trained using ImageNet-100 dataset. (a) Supervised learning — as we include more classes in the training that starts from scratch, both $\mathsf{trank}$ and $\mathsf{erank}$ increase. Algebraic rank remains at the maximum value. (b) Unsupervised learning with SimCLR loss — as the unsupervised representation learning proceeds, both $\mathsf{trank}$ and $\mathsf{erank}$ increase. Algebraic rank remains at the maximum value.
@@ 4 Experiments
We provide an overview of our experimental settings in Section 4.1. Subsequently, we demonstrate the efficacy of our method for enhancing novel task performance in Section 4.2 and for mitigating catastrophic forgetting in Section 4.3. In Section 4.4, we demonstrate that our method can improve performance of eleven well-known backward-compatible methods. In Section 4.5, further analysis is provided. In Section 4.6, an ablation study is presented. In Section 4.2 and Section 4.3, we focus our analysis on the UCIR method, given its prevalent adoption across recent state-of-the-art methodologies. Code is available at: https://github.com/Wonseok1017/NN-RFR
(a) S=10
(b) S=5
(c) S=2
@@
4.1 Settings
Datasets:
We employ CIFAR-100 and ImageNet-100, two widely adopted benchmark datasets in CIL. To maintain consistency with prior studies, we follow the standard class orderings proposed in (Rebuffi et al., 2017) for all evaluations except for the evaluation of PODNet (Douillard et al., 2020) and IL2A (Zhu et al., 2021), for which we utilize the class orderings defined in PODNet and IL2A, respectively. To evaluate the CIL methods, all classes in datasets are divided into multiple tasks. The initial 50 classes are designated for the base task, while the remaining classes are split into novel tasks, where the size of each split is either 10, 5, or 2. In this study, we denote split sizes of 10, 5, and 2 as S=10, S=5, and S=2, respectively. To pre-process the datasets, we sequentially apply two simple augmentations: random cropping followed by resizing to the original dimensions (32x32 for the CIFAR dataset and 224x224 for the ImageNet dataset). Additionally, there is a 50% probability of applying random horizontal flipping. Afterward, we normalize the color channels by subtracting the mean values ((0.4914, 0.4822, 0.4465) for CIFAR and (0.485, 0.456, 0.406) for ImageNet) and dividing by the respective standard deviations ((0.247, 0.243, 0.261) for CIFAR and (0.229, 0.224, 0.225) for ImageNet).
Implementation Details:
In this study, ResNet-18 (He et al., 2016) is employed as the base model to investigate. To reproduce the results of iCaRL (Rebuffi et al., 2017), LwF (Li and Hoiem, 2017), SI (Zenke et al., 2017), EEIL (Castro et al., 2018), MAS (Aljundi et al., 2018), RWalk (Chaudhry et al., 2018), BiC (Wu et al., 2019), IL2M (Belouadah and Popescu, 2019), and UCIR (Hou et al., 2019), we utilize the open-source codebase provided by FACIL (Masana et al., 2022) and retain their original hyperparameters unchanged. For PODNet (Douillard et al., 2020), IL2A (Zhu et al., 2021), and AFC (Kang et al., 2022), we employ their own open-source codebases and retain their original hyperparameters unchanged. To incorporate RFR into the aforementioned methods, we only modify their respective loss functions during the training of the base session, as specified by Eq. (6). To determine the optimal RFR loss coefficient, $\alpha$, we performed a grid search over $\alpha\in\{0.05,0.10,0.15,0.20\}$. Meanwhile, the remaining configurations, such as the loss functions applied during the training of novel sessions and the default hyper-parameter settings, remain unchanged. To evaluate the performance of the models, we adopt the standard metric of average incremental accuracy (AIC) proposed in (Rebuffi et al., 2017).
4.2 Superiority in forward compatibility
4.3 Mitigation in catastrophic forgetting
@@
To analyze improvement in catastrophic forgetting, we conduct three investigations. First, we examine the influence of our method on the weight changes of the feature extractor. As novel sessions progress, the $L_{2}$-weight distance between the feature extractor learned in the base session and that in the novel session increases, as demonstrated in Figure 10. However, with the integration of RFR, the weight distances are significantly reduced across all sessions. The average reductions in weight distances are 8.70, 8.13, and 6.79 for split sizes of 10, 5, and 2, respectively. Additionally, Figure 10 reveals that the weight distance with the feature extractor from the immediate previous session also increases less when our method is employed, showing average reduction of 6.55, 4.95, and 2.99 for split sizes of 10, 5, and 2, respectively.
@@ 4.4 Consistent improvements over eleven existing methods
iCaRL (Rebuffi et al., 2017) 49.83±2.33 46.86±2.38 44.69±1.97 53.77±1.68 48.41±8.70 49.78±3.65
    with RFR 50.56±1.47 48.48±1.17 44.99±0.97 57.17±1.66 55.62±0.71 54.40±1.63
@@
EEIL (Castro et al., 2018) 37.81±2.55 24.48±1.18 32.37±2.52 52.04±1.90 48.65±2.34 43.95±1.65
    with RFR 39.52±0.38 25.79±1.18 33.46±0.69 54.74±0.91 51.64±0.80 47.89±0.60
@@
RWalk (Chaudhry et al., 2018) 35.75±1.63 23.34±2.06 27.02±5.36 41.08±1.91 21.44±5.69 20.81±2.12
    with RFR 39.00±1.15 24.44±1.06 32.85±1.40 44.83±1.87 35.42±1.71 22.68±4.70
@@
IL2M (Belouadah and Popescu, 2019) 40.85±2.22 24.92±1.12 33.41±2.12 56.78±1.34 52.35±2.16 47.10±1.85
    with RFR 42.54±0.49 25.19±0.41 35.20±0.78 58.40±1.04 55.25±0.78 50.09±0.83
@@
PODNet(NME) (Douillard et al., 2020) 68.47±1.27 67.09±1.19 65.16±1.02 60.42±0.53 50.93±0.73 36.36±0.30
    with RFR 69.51±0.56 67.66±0.63 65.59±0.48 64.08±0.45 56.34±0.75 41.03±2.15
@@
To demonstrate the efficacy of RFR in enhancing performance, we conduct comprehensive experiments over eleven well-known existing works, all utilizing a single backbone. Notably, dynamic network methods (Yan et al., 2021; Douillard et al., 2022; Wang et al., 2022b, a) were excluded, as they generally use separate backbones for each task. This design inherently makes them more resistant to catastrophic forgetting (Zhou et al., 2024), thereby reducing their reliance on forward compatibility. However, this advantage comes at the cost of increased computational demands with each new session.
The results presented in Table 1 and Table 2 provide compelling evidence of the significant and consistent performance improvements achieved through the integration of RFR into the previous works. Specifically, we observe notable average performance improvements for the state-of-the-art works, including a 3.74% average improvement for UCIR, a 2.63% average improvement for PODNet (NME), a 1.24% average improvement for PODNet (CNN), and a 0.36% average improvement for AFC. Particularly, a remarkable performance improvement is observed in the case of UCIR for CIFAR-100 and its further analysis is shown in Figure 12. The achieved improvements are 3.14% increase for S=10, 5.59% for S=5, and 8.49% for S=2.
@@
ResNet-50 UCIR 67.10±0.20 61.42±0.61 51.62±0.14
    with RFR 68.57±0.12 64.81±0.15 57.71±0.31
@@
ViT-Small UCIR 50.45±0.35 49.03±0.27 46.24±0.79
    with RFR 51.86±0.35 50.50±0.47 47.96±1.03
@@ 4.5 Further analysis
In addition to evaluating RFR’s performance under the widely used benchmark settings, we extend our evaluation to a larger-scale dataset. Due to the high computational demands of the ImageNet-1000 dataset and our limited resources, we evaluate RFR using only a single method. As shown in Table 4, RFR continues to demonstrate its effectiveness for ImageNet-1000 dataset. Furthermore, we analyze RFR’s performance with larger-scale models. The results presented in Table 5 underscore RFR’s efficacy across a range of larger-scale models, including ResNet-50 and ViT-Small (Lee et al., 2021).
@@ 4.6 Ablation study
B=10 60.31±0.22 60.41±0.34 +0.10
B=20 63.16±0.34 65.11±0.29 +1.95
B=30 65.17±0.25 66.98±0.21 +1.81
B=40 67.22±0.25 69.53±0.12 +2.31
B=50 66.30±0.36 69.45±0.29 +3.15
B=60 70.44±0.13 73.06±0.36 +2.62
B=70 71.57±0.39 74.20±0.20 +2.63
B=80 73.07±0.18 75.95±0.09 +2.88
B=90 73.41±0.43 75.98±0.19 +2.57
Time cost analysis:
We conducted a time cost analysis of Base Session Training using various datasets. As demonstrated in Table 7, the experimental results indicate a slight increase in the average time; however, considering the standard deviation, this difference appears to be negligible. This result is likely due to the highly efficient implementation of Singular Value Decomposition (SVD) in the CUBLAS library (Lahabar and Narayanan, 2009), which is the default linear algebra library used by PyTorch.
Table 7: Time Cost Analysis of Base Session Training with Various Datasets. The average training time (in seconds) per epoch was measured for each dataset. Each model was trained for three runs using an NVIDIA RTX 3090 GPU.
CIFAR100 ImageNet-100 ImageNet-1000
B=50 B=50 B=500
UCIR 20.22±0.18 145.00±0.76 995.53±15.10
    with RFR 20.50±0.21 146.48±3.20 996.96±15.30
@@ 5.1 Comparison with CwD
Baseline 66.30±0.36 60.57±0.56 52.74±0.72 66.11±0.63 63.49±0.45 59.61±0.53
    with CwD 67.06±0.12 62.71±0.36 56.39±0.30 66.01±0.75 63.63±1.02 60.14±1.16
    with RFR 69.45±0.29 66.16±0.10 61.23±0.13 67.13±0.86 64.72±1.01 61.49±0.71
@@
5.2 Comparisons to the most recent methods
Table 9: Comparisons to the most recent CIL methods.
Method CIFAR100 (B=50) ImageNet-100 (B=50)
S=10 S=5 S=10 S=5
DCMI (Qiu et al., 2024) 67.90 66.80 70.50 70.00
FCS (Li et al., 2024) 62.13 60.39 N/A 61.76
MRFA (Zheng et al., 2024) 74.73 N/A 79.29 N/A
AFC (Kang et al., 2022) with RFR 70.85±0.21 67.82±0.37 77.19±0.08 75.47±0.23
While the benchmark methods presented in Section 4.4 are widely adopted in CIL due to their popularity and reliably reproducible performance, they do not include the most recent methods in CIL. In this section, we compare the performance of RFR with the most recent methods (Qiu et al., 2024; Li et al., 2024; Zheng et al., 2024). Since we could not reproduce the originally reported performance of these methods, we present their reported performance from their respective papers and compare it with our performance (AFC with RFR). The results are shown in Table 9. Although MRFA (Zheng et al., 2024) demonstrates the best performance when S=10, our method achieves the second-best performance for S=10. For S=5, our method surpasses the other recent methods. These results indicate that the benchmark performance comparisons in Section 4.4 remain competitive also for the recent methods.
@@ 5.3 Increasing representation rank during novel sessions
RFR increases effective rank during the base session only. It is also possible to increase effective rank during novel sessions with the goal of acquiring additional features that may be useful in the subsequent novel sessions. Such a strategy, however, can also intensify catastrophic forgetting during novel sessions. When additional regularization, such as increasing the effective rank, is applied during the training of a novel task, it results in more significant changes to the model compared to when only the novel task is learned. This occurs because the model must simultaneously learn both the novel task and the regularization task. Consequently, the regularization intended to enhance the model’s forward compatibility with subsequent tasks introduces a trade-off, increasing catastrophic forgetting of previous tasks. To investigate the overall effect, we have conducted an experiment and the results are shown in Table 10. While RFR demonstrates effectiveness in enhancing feature richness during the base session, its influence on overall performance is marginal when also applied to novel sessions. Therefore, we have chosen to apply effective rank regularization only during the base session.
@@
BiC 58.06±0.73 58.29±0.38 +0.23 48.02±2.33 46.60±1.91 -1.41 27.74±3.11 27.64±1.83 -0.10
EEIL 39.52±0.38 39.44±0.63 -0.08 25.79±1.18 25.62±1.10 -0.17 33.46±0.69 33.16±1.07 -0.29
iCaRL 50.56±1.47 50.50±2.04 -0.06 48.48±1.17 48.23±1.14 -0.25 44.99±0.97 44.81±1.18 -0.18
IL2M 42.54±0.49 42.76±0.83 +0.21 25.19±0.41 26.22±1.77 +1.03 35.20±0.78 34.87±0.67 -0.32
LwF 40.98±0.34 40.69±0.81 -0.28 28.82±4.17 26.70±0.39 -2.12 34.46±0.63 34.63±0.75 +0.17
MAS 39.54±0.71 39.77±0.66 +0.23 28.80±4.52 29.03±3.50 +0.23 33.36±0.52 33.91±0.91 +0.55
SI 39.41±0.60 39.51±0.58 +0.10 24.02±0.38 23.89±0.19 -0.13 33.48±0.80 33.60±0.52 +0.12
RWalk 39.00±1.15 38.91±1.39 -0.09 24.44±1.06 25.96±3.06 +1.53 32.85±1.40 33.04±1.20 +0.19
UCIR 69.45±0.29 69.57±0.23 +0.12 66.16±0.10 66.55±0.16 +0.39 61.23±0.13 61.27±0.29 +0.04
@@
5.4 Practical application perspective
In this study, we demonstrated that RFR is effective across models of various sizes and architectures, including both convolution-based architectures (e.g., ResNet-18 and ResNet-50) and transformer-based architectures (e.g., ViT-Small). Given that RFR introduces minimal additional training time and is effective in single-network methods, which are relatively resource-efficient (Zhou et al., 2024), it suggests that RFR may offer greater advantages in environments with limited training resources, such as edge devices.
5.5 Limitations
The effectiveness of RFR depends on the accurate estimation of the effective rank of the representation. Since the effective rank is mathematically upper-bounded by the batch size, RFR regularization tends to be less efficacious with smaller batch sizes. Empirical evidence indicates that batch sizes greater than 32 are sufficient to achieve consistent and reliable performance.
We demonstrate the efficacy of RFR in enhancing forward compatibility by applying it to CIL methods that use a single backbone. The limited effectiveness of RFR when applied to dynamic network methods with increasing backbones underscores a limitation of this approach. Additionally, due to the high computational demands of the ImageNet-1000 dataset and limited resources, we evaluated RFR using only a single method.
We have followed previous works and presented the main results in Table 1 and Table 2. Additionally, Table 3 focuses on non-exemplar based CIL, Table 4 addresses large-scale dataset, and Table 5 covers large-scale models. To ensure the generalizability of the proposed method, it would also be beneficial to consider real-world scenarios across a variety of application domains.